A researcher is interested in finding a 95% confidence interval for the mean number minutes students are concentrating on their professor during a one hour statistics lecture. The study included 121 students who averaged 37.2 minutes concentrating on their professor during the hour lecture. The standard deviation was 13.3 minutes. Round answers to 3 decimal places where possible.
a. To compute the confidence interval use a
?
distribution.
b. With 95% confidence the population mean minutes of concentration is between
and
minutes.
c. If many groups of 121 randomly selected members are studied, then a different confidence interval would be produced from each group. About
percent of these confidence intervals will contain the true population mean minutes of concentration and about
percent will not contain the true population mean minutes of concentration.

Answer by Boreal(15235)

I would use a t-distribution because the sd used is from the sample. Some might use a z- because the number is large. There won't be a big difference but t is more accurate
half-interval is t(0.995,df=120)*ss/sqrt(n)=2.617*13.3/11
=3.164 min
the interval is (34.036, 40.364)units min.
-
95% for first part
5% for second part